OpenStudy (anonymous):

Twenty-two is what percent of 55? A. 50% B. 40% C. 22% D. 250%

OpenStudy (anonymous):

Use the formula: Remeber: X = is ___ ___ 100 of

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

55 * (x)=22 x = 22/55 = 0.4 = 40%

OpenStudy (anonymous):

thanks